OpenStudy (anonymous):

im really confused and tired can someone help me?Estimate the roots of the equation by graphing. Round to the nearest tenth and write in set form, such as {-1.2,3.4}. x2 - 6= 0

OpenStudy (anonymous):

how do i solve that?

Parth (parthkohli):

add 6 to both sides, take the square root.

OpenStudy (anonymous):

x^2=6?

Parth (parthkohli):

Yes, now what are the square roots of 6?

OpenStudy (anonymous):

2.44948974278

Parth (parthkohli):

Now round to the nearest tenth

OpenStudy (anonymous):

2.4

Parth (parthkohli):

Note that there is a negative sqrt too.

OpenStudy (mathstudent55):

An equation of the form x^2 = a, a > 0 has two roots: x = sqrt(b) and x = -sqrt(b) Since you have x^2 = 6, just find the square root of 6. Then one root is the positive square root of 6 and the other root is the negative square root of 6.

OpenStudy (anonymous):

2.4,-2.4

Parth (parthkohli):

that is right!

OpenStudy (anonymous):

thats the answer?

Parth (parthkohli):

Yes.

OpenStudy (koikkara):

well.....is it 3??